# Learned Kernels for Sparse, Interpretable, and Efficient Medical Time Series Processing

### v3

@@ 4.4 Correlated Kernel Pruning
When pruning a kernel, we need to update the weight and biases of the remaining kernel to correct for the removed kernel’s contribution. Without loss of generality, assume that $\text{eff}_{j}>\text{eff}_{k}$. We update the weight and bias of the remaining kernel $i$ as follows:
@@
where $b_{j}$ and $b_{k}$ are the biases for weight $w_{j}$ and $w_{k}$ respectively. In essence, the bias of the two kernels are simply combined into one term to compensate for pruning one kernel, and the weight of the remaining kernel is increased proportionally to account for the loss in signal from the pruned kernel. This technique allows for the removal of 22% of the parameters of our largest model, with less than a 2% absolute impact on performance on all datasets. Moreover, this technique is adaptable, allowing for the removal of more parameters at a greater performance cost, or a smaller number of parameters at a lesser performance cost. Notably, the choice of Euclidean distance is arbitrary, and any distance metric can be chosen for the pruning step. Indeed, one can use a variety of distance metrics with similar results (see Appendix Table 7).
5 Experimental Setup
5.1 PPG Segmentation
5.2 Training
5.3 Evaluation
@@
5.4 Baselines
We compare to the same baselines tested in Guo et al. [1]. Namely, a convolutional neural network sliding window approach, a template-matching approach, and a ResNet-34-based classifier with segmentation performed via GradCAM [42] or SHAP [43]. The baselines are discussed in more detail in Appendix A.
@@ 5.5.3 Baselines
We compare against a common deep learning architecture for medical signal processing, namely a 1D-ResNet. The baseline is discussed in Appendix A.
5.6 Interpretability
We trained another learned kernel classifier to distinguish between $1^{\circ}$ atrioventricular block (1140 samples) and sinus rhythm (8125 samples) to demonstrate the interpretability of our method. We chose $1^{\circ}$ atrioventricular block to demonstrate the interpretability of our method as it is defined by a single salient feature (the duration between the P-wave and QRS-complex), rather than a set of global features (e.g., irregularly irregular beats in atrial fibrillation). We gather this data from the Zheng et al. dataset.
@@ 6.1.1 Test Set Performance
We find that our largest model significantly exceeds the performance (measured in DICE score) of all baselines except the current state-of-the-art models (Segade and Tiny-PPG), for which it achieves $>$99% of their DICE-score on DaLiA and WESAD, with less than 2% the parameter count of Segade and half the parameters of Tiny-PPG. Our largest model even manages to significantly outperform all other baselines on the most challenging and least in-distribution dataset (TROIKA). Our medium model, at 0.4% of the parameter count of Segade, achieves at least 98% of Segade’s DICE score on all datasets and matches Segade’s performance on TROIKA. Lastly, our smallest model, at 0.06% of the parameter count of Segade, reaches 94% of the DICE score of Segade on DaLiA and WESAD, though struggles on TROIKA, reaching 86% of the DICE score of Segade. Detailed results and parameter counts are in Table 1. Furthermore, we find that our model is robust to noisy inputs, showing little degradation in DICE score when varying levels Gaussian noise are added to the test signal, even without augmenting training data with Gaussian noise (see Appendix Figure 9).
@@
Table 1: Various baselines vs. our approach at varying parameter counts. Parameter counts above 1M are in red and parameters below 50K are in blue. Italicized results are at least 99% of the performance of the best result. Our methods are in bold.
6.1.2 Scaling Laws
In this section, we explore the scaling behavior of our approach by gradually increasing the number of kernels and evaluating the performance on the test set. Our main objective is to understand the relationship between the parameter count and test-set performance, which will help us identify the potential limits and robustness of our approach.
We note that it is possible that a theoretical performance limit exists due to mislabeling in the test set; this is discussed further in Appendix section 5.
@@ 6.1.3 Interpretability
where $\|\mathbf{k}_{m}\|_{2}^{2}=\sum_{j}k^{2}_{m,j}$ is the sum of each squared component of the $m^{th}$ kernel, $\textit{b}_{m}$ is the $m^{th}$ kernel’s bias, and $\textit{w}_{m}$ is the $m^{th}$ kernel’s weight. This, by definition, computes the value added to the final signal by that kernel when the kernel perfectly overlaps with an exact template match. Furthermore, we also note that in the case of a normalized kernel and bias (i.e. $\sum_{j}\textbf{k}^{2}_{m,j}+b_{m}=1$), the kernel importance directly equals the weighting of that kernel. We run this kernel analysis across all 10 cross-validation models for the small, medium, and large model sizes.
We also compute and record the output values of each kernel group over the whole test dataset. Essentially, we perform the convolution and weighting operation for each set of kernels grouped by size (long, moderate, or short), along the entire test dataset, concatenating the output values to one array. We further segregate the output values depending on whether they were generated by PPG signals labeled as artifact or clean. In the framework of our model, this is equivalent to computing:
@@
We notice a clear pattern between kernel size and kernel importance. We find that for the medium and large models, the long kernels have significantly more positive ($p<10^{-13}$) mean importance than the short and moderate kernels, with significance computed via two-tailed independent samples t-tests. For the small model, the long and moderate kernels have a mean importance significantly more positive than the small kernels ($p<10^{-6}$) (see Figure 4). These results indicate that long kernels generally learn to recognize poor-quality signal features, while medium and short kernels learn to recognize clean PPG features. We investigate this claim empirically by computing the activations of each kernel size group over the entire test dataset, grouping activation values by whether they were generated by artifact-containing PPG signals or clean signals. We find the claims generated by our mathematical analysis of kernel importance to be consistent with our empirical findings, with long kernels producing positive values most of the time, moderate kernels producing mostly negative or mixed positive/negative values (as seen in the medium-sized model), and short kernels producing mostly negative values, as shown in Figure 4. We demonstrate one such 30-second example of this in Figure 5. The agreement between our mathematical exploration of kernel importance and our empirical exploration of the values produced by various kernel groups over the test sets demonstrates that we can draw meaningful conclusions of our model without relying on post hoc explanability methods. These results are shown in Figure 4 by comparing the top and bottom rows.
Figure 4: The computed “kernel importance” of each kind of kernel in each model size (top row) and the empirical average activation of each kernel group observed over the clean and artifact parts of the dataset (bottom row). Significant differences in computed “kernel importances” are reflected in the empirical output values of those kernel groups over the test set. (*) denotes statistical significance at p $<10^{-6}$ via a two-tailed independent samples t-test with Bonferroni correction. All differences are significant between the empirical kernel output due to the extremely large sample size.
Next, we note that because our smallest model consists of only 12 kernels, we can easily visually inspect each kernel. They are all shown in Appendix Figure 8. This is in stark contrast to compared to most deep-learning models, where the parameters are typically so numerous that it is intractable to manually inspect each parameter; here, all of the kernels can be shown on one page. Furthermore, parameters in deep neural networks are typically uninterpretable, whereas our model represents a template-matching paradigm with a waveform.
@@ 6.1.4 Pruning Experiments
We find that applying our pruning procedure on long kernels provides the greatest parameter reduction with the least performance trade-off. Pruning moderate and/or short kernels impacts performance more heavily and ultimately does not prune as many parameters as pruning long kernels, since each long kernel contains more parameters by definition. Ultimately, the choice of prune-to-performance ratio is up to the user and specific use case, though we provide example pruning schemes in Appendix Table 6 which attempt to maintain performance around 96% of the mean original performance. Notably, as the model size increases, we are able to remove progressively more parameters with less of a decrease in performance. We hypothesize this may be due to increased capacity of larger models and perhaps greater learned redundancy.
@@ 6.2 Atrial Fibrillation Detection
We tested our 12-kernel, 72-kernel, and 384-kernel models against a 16-layer ResNet. Despite the ResNet significantly outperforming all of our learned kernel models on the train set cross-validation metrics (Table 2), we find that our models, having several orders of magnitude fewer parameters, generalize significantly better to the unseen holdout dataset, shown in Table 3.
We also find that our models train and generalize significantly better in the low-data regime. Training on only 1% of the dataset, our method achieves higher F1 and AUC-ROC scores than ResNet-16 on both the train and holdout set. We hypothesized this could be due to overparameterization of our ResNet-16 model, thus as a control, we train a simple four-layer convolutional neural network with nearly the same parameter count as our large learned kernel model. We find that even when controlling for parameter count, deep neural networks struggle to learn in the low data regime, achieving near-random performance, while our method maintains a stable performance across parameter counts (Table 4).
@@
Table 2: Metrics on the train set, computed via 10-fold cross-validation.
@@ 6.2.1 Interpretability
To demonstrate the built-in interpretability of our method for classification tasks, we train a large learned kernel model to distinguish between $1^{\circ}$ atrioventricular block and sinus rhythm. We then apply Algorithm 1 to create a contribution map for each class. We find that our importance maps reliably identify the PR-interval (the defining factor) as the most contributory factor to the classification of $1^{\circ}$ atrioventricular block, despite never explicitly being given this signal (Figure 6). Importantly, this is not a post-hoc methodology (such as SHAP or GradCAM) that seeks to explain a model’s behavior but is not guaranteed to be faithful. Instead, this method numerically computes the precise contribution of each part of the signal to the classification.
Figure 6: Various examples of $1^{\circ}$ atrioventricular block from the Zheng et al. dataset, along with the corresponding contribution maps as computed by algorithm 1. Redder colors indicate greater contribution to the $1^{\circ}$ atrioventricular block, while bluer colors indicate lesser contribution.
@@ 7 Discussion
In this study, we proposed a novel architecture for medical signal processing, which maintains performance on par or exceeding deep learning methods, all while remaining interpretable and efficient. We demonstrated the efficacy of our approach using PPG artifact detection as a case study as well as single-lead atrial fibrillation detection. For PPG artifact detection, our approach achieved near state-of-the-art performance on three benchmark datasets with several orders of magnitude fewer parameters than deep learning methods and showed better generalization to out-of-distribution data, as evident by the markedly improved performance in the most out-of-distribution task, the TROIKA dataset. This method’s simplicity and interpretability allow for the inspection of the model’s inner workings, providing insights into the contribution of each kernel to the overall signal quality assessment. Furthermore, our method demonstrates extremely low compute and memory demands, setting it apart from deep-learning-based approaches. This enables high-quality signal assessment on low-power wearable devices.
@@
Our results demonstrated that SMoLK can effectively capture relevant features for identifying good and poor-quality PPG signals, as well as provide a directly interpretable, lightweight, and performant classifier. The scaling behavior of our approach showed a power-law relationship between parameter count and test-set performance. Our analysis of the kernel importance revealed that larger kernels primarily contribute to the identification of poor-quality signal features, while smaller and medium-sized kernels capture normal PPG features. Furthermore, we demonstrate that our single-lead classifier is directly interpretable, correctly identifying defining features of arrhythmias. Importantly, our work demonstrates that for some complex tasks, simple and interpretable models can perform just as well as deep neural networks and are even preferable.
Furthermore, we developed novel parameter reduction techniques to improve the compute-to-performance ratio, keeping in mind the low-power limitations of wearable devices. We demonstrated that we can remove over 22% of parameters in large models and maintain above 98% of the original performance on a PPG segmentation task.
@@
We also demonstrated that, owing to the simplicity and near linearity of our model, we can naïvely quantize the model weights with no impact on performance. This is in stark contrast to deep neural networks, where the chaotic properties of the system can amplify and propagate error due to quantization, leading to performance degradation. Quantization is also often a non-trivial process in deep neural networks as a result of dynamic range issues, again owing to the chaotic nature of neural networks. Our proposed method suffers from none of the aforementioned complexities.
@@
All in all, our largest model can fit in under 100 KB of memory when cast to float16 and can run in under 75 KB after weight pruning. Our smallest model runs using under 3 KB of memory cast to float16. Computing a raw (i.e., without post-process smoothing) signal score is extremely computationally efficient, with our largest model requiring on the order of 10 MFlops/second of signal processing. This memory-efficient and compute-efficient approach means that designers of low-power, wearable devices can now achieve state-of-the-art PPG signal quality assessment on extremely stringent compute budgets. Our models are so compute-efficient that they could feasibly be implemented on microcontrollers, for example, on consumer pulse oximeters. This starkly contrasts with deep learning methods which are memory and compute-intensive. Additionally, implementation of our architecture is simple and straightforward to adapt to lower-level programming languages, whereas deep learning-based methods require the time-intensive re-implementation of complex deep learning operations into lower-level devices. Furthermore, for segmentation tasks, SMoLK’s output only has local dependencies on the input signal. In other words, one does not need to wait for chunks of signal to be acquired before running the segmentation algorithm (as is the case for deep networks, which often have global dependencies and a fixed input dimensionality), nor does one need to stitch together multiple prediction windows. Instead, the output of the segmentation algorithm can be computed in real-time as the signal is acquired.
We also highlighted potential limitations in the datasets used for evaluation, as some labeling inaccuracies and discrepancies were observed in the test set samples with the poorest performance. This suggests that the practical maximum achievable performance for PPG artifact detection on these test sets may be limited by the quality of the available ground truth labels.
@@
In conclusion, our method provides a robust, interpretable, and efficient approach for PPG signal quality assessment, making it competitive with deep learning approaches in performance and yet suitable for deployment on low-power devices.
@@
While our findings are promising, further research is needed to validate our approach in other medical signal-processing applications. Nonetheless, our study provides a solid foundation for the development of efficient, interpretable, and high-performing medical signal processing models. The application of our method to other physiological signals, such as electrocardiography, and tasks could be explored, as well as the integration of our approach with other wearable devices like consumer pulse-oximeters, smart-rings, smart watches, and other low-profile devices for real-time signal quality assessment. Our work paves the way for the integration of deep-learned quality signal processing into low-power wearable devices. This could significantly advance the field of wearable health technology, enabling continuous, real-time monitoring of biomarkers with high accuracy and minimal computational resources.

### v4

@@ 4.2.2 Pruning
When pruning a kernel, we need to update the weight and biases of the remaining kernel to correct for the removed kernel’s contribution. Without loss of generality, assume that $\text{eff}_{j}>\text{eff}_{k}$. We update the weight and bias of the remaining kernel $j$ as follows:
@@
where $b_{j}$ and $b_{k}$ are the biases for weight $w_{j}$ and $w_{k}$ respectively. In essence, the bias of the two kernels are simply combined into one term to compensate for pruning one kernel, and the weight of the remaining kernel is increased proportionally to account for the loss in signal from the pruned kernel. This technique allows for the removal of 22% of the parameters of our largest model, with less than a 2% absolute impact on performance on all datasets (Appendix Table 7). Moreover, this technique is adaptable, allowing for the removal of more parameters at a greater performance cost, or a smaller number of parameters at a lesser performance cost. Notably, the choice of Euclidean distance is arbitrary, and any distance metric can be chosen for the pruning step. Indeed, one can use a variety of distance metrics with similar results.
We provide three example pruning results at various model scales, with a target performance decrease of no greater than 5% (Appendix Table 6). We would like to note that one can experiment with this pruning process via our open-source GitHub repository.
Furthermore, we show pruning results using various distance metrics perform relatively consistently, indicating a robustness to the chosen distance metric. We test three additional distance metrics on our large model: cosine similarity between power spectra (cosine distance in frequency space), cosine similarity in feature space (the cosine similarity of the feature vectors themselves), and Manhattan distance (Appendix Table 7).
4.2.3 Evaluation
@@
4.2.4 Baselines
We compare to the same baselines tested in Guo et al. [41]. Namely, a convolutional neural network sliding window approach, a template-matching approach, and a ResNet-34-based classifier with segmentation performed via GradCAM [42] or SHAP [43]. Further details are given in the Baseline Methods section.
@@ 4.3.3 Baselines
We compare against a common deep learning architecture for medical signal processing, namely a 1D-ResNet, as well as several previously state-of-the-art low-parameter count convolutional neural networks. These baselines are outlined in the Baseline Methods section.
4.3.4 Interpretability
We trained another learned kernel classifier to distinguish between $1^{\circ}$ atrioventricular block (1140 samples) and sinus rhythm (8125 samples) to demonstrate the interpretability of our method. We chose $1^{\circ}$ atrioventricular block to demonstrate the interpretability of our method as it is defined by a single salient feature (the PR-interval), rather than a set of global features (e.g., irregularly irregular beats in atrial fibrillation). We gather this data from the Zheng et al. dataset.
4.4 Baseline Methods
4.4.1 Convolutional Sliding Window
4.4.2 Template-Matching
We follow the same approach as Guo et al. [41], which was inspired by Lim et al.[60]. Signals were first divided from the PPG-DaLiA training set into separate pulses via peak detection. 10 clean (artifact-free) pulses were then selected to serve as our standard templates for comparison with test pulses. Each pulse from the test data was compared to these 10 templates, using the dynamic time warping (DTW) distance metric to measure similarity. The smallest DTW distance (out of the 10 comparisons with the templates) was then identified. The range of the DTW distance function is $[0,\infty)$, but empirically we find that the DTW distances between our templates and the PPG signals fall mostly within the range $[0,10]$. We created a binary predicted label by labeling a segment as an artifact if the minimum DTW distance is at least 1. Other samples were given “non-artifact” predicted labels. The threshold of DTW was chosen by testing thresholds between 0 and 10 over the entire train set and choosing the threshold $\beta$ that produced the highest DICE score (which was 1). To summarize, if the minimum DTW distance $\alpha$ exceeds our threshold $\beta$, we classify the entire pulse (and all its timesteps) as an artifact.
4.4.3 GradCAM & SHAP
In this baseline experiment, we compare to the results achieved by Guo et al. [41]. Briefly, Guo et al. conducted their experiment by employing the Resnet-34 architecture proposed by Dai et al (2016) for 1D binary classification (identifying ‘clean’ or ‘artifact’). This architecture, traditionally used for image classification, was repurposed for time series analysis via GradCAM [42]. Considering the relatively small size of the training dataset, they employed transfer learning on a pre-trained Resnet34-1D PPG signal quality classifier, as described by Zhang et al (2021) [61]. This pre-trained model was originally trained on a UCSF PPG dataset acquired by Zhang et al [61] from 3764 intensive-care unit patients according to Pereira et al (2019) [4]. The last two residual blocks, global average pooling, and the final dense layer were subsequently retrained by Guo et al. To train a classification model, they created ground truth labels such that any signal that contained artifact timesteps was labeled as an artifact, while all other signals were labeled as non-artifacts. After this labeling process, the training set consisted of 175 clean signals and 3261 artifact signals.
While the Resnet34-1D network is fundamentally a classifier, the goal is instead to obtain segmentation labels. Assuming that the model would focus on artifacts when tasked with predicting an artifact signal, Guo et al. employed two post-hoc explanatory methods to generate segmentation masks: SHAP [43] and GradCAM [42].
4.4.4 Segade
4.4.5 Tiny-PPG
4.4.6 ResNet-16
4.4.7 Jia-CNN
We use the same three convolutional neural network architectures as Jia et al. [63] (referred to as Jia-CNN), though we change the number of final output neurons from two to three to accommodate the difference in class number. Additionally, Jia-CNN was designed to receive an input signal at 250 Hz, thus we change only the resampling frequency of our preprocessing pipeline, leaving all else the same for an equal comparison. We follow the same training paradigm as ResNet-16, though we notably raise the learning rate to $0.005$ for Jia-CNN1 and Jia-CNN3. We set the learning rate to $0.002$ for Jia-CNN2, as we find that on some train splits, the model is unable to converge with higher learning rates.
@@ 5.1.1 Test Set Performance
We conduct our tests on several PPG datasets, DaLiA [46], WESAD [47], and TROIKA [48], against several deep learning (Segade [41] and Tiny-PPG [36]) and traditional machine learning methods. A full description of the datasets, pre-processing, and baseline experiments is given in the Methods section. We find that our largest model significantly exceeds the performance (measured in DICE score) of all baselines except the current state-of-the-art models (Segade and Tiny-PPG), for which it achieves $>$99% of their DICE-score on DaLiA and WESAD, with less than 2% the parameter count of Segade and half the parameters of Tiny-PPG. Our largest model even manages to significantly outperform all other baselines on the most challenging and least in-distribution dataset (TROIKA). Our medium model, at 0.4% of the parameter count of Segade, achieves at least 98% of Segade’s DICE score on all datasets and matches Segade’s performance on TROIKA. Lastly, our smallest model, at 0.06% of the parameter count of Segade, reaches 94% of the DICE score of Segade on DaLiA and WESAD, though struggles on TROIKA, reaching 86% of the DICE score of Segade. Detailed results and parameter counts are in Table 1. Furthermore, we find that our model is robust to noisy inputs, showing little degradation in DICE score when varying levels Gaussian noise are added to the test signal, even without augmenting training data with Gaussian noise (see Appendix Figure 9). We also find that it is empirically challenging to overfit on our train set even at high parameter counts and that there may be a potential performance ceiling due to data labeling inaccuracies (see Appendix B and C).
@@
Table 1: PPG Segmentation Performance. Various baselines vs. our approach at varying parameter counts on the PPG segmentation task, measured by DICE score. Italicized results are at least 99% of the performance of the best result. Our methods are in bold.
@@ 5.1.2 Interpretability
where $\|\mathbf{k}_{m}\|_{2}^{2}=\sum_{j}k^{2}_{m,j}$ is the sum of each squared component of the $m^{th}$ kernel, $\textit{b}_{m}$ is the $m^{th}$ kernel’s bias, and $\textit{w}_{m}$ is the $m^{th}$ kernel’s weight. This, by definition, computes the value added to the final signal by that kernel when the kernel perfectly overlaps with an exact template match. Furthermore, we also note that in the case of a normalized kernel and bias (i.e., $\sum_{j}\textbf{k}^{2}_{m,j}+b_{m}=1$), the kernel importance directly equals the weighting of that kernel. We run this kernel analysis across all 10 cross-validation models for the small, medium, and large model sizes.
We notice a clear pattern between kernel size and kernel importance. We find that for the medium and large models, the long kernels have significantly more positive ($p<10^{-13}$) mean importance than the short and moderate kernels, with significance computed via two-tailed independent samples t-tests. For the small model, the long and moderate kernels have a mean importance significantly more positive than the small kernels ($p<10^{-6}$) (see Figure 3). These results indicate that long kernels generally learn to recognize poor-quality signal features, while medium and short kernels learn to recognize clean PPG features.
Figure 3: Kernel Statistics. The computed “kernel importance” of each kind of kernel in each model size (top row) and the empirical average activation of each kernel group observed over the clean and artifact parts of the dataset (bottom row). Significant differences in computed “kernel importances” are reflected in the empirical output values of those kernel groups over the test set. Significance was computed via a two-tailed independent samples t-test with Bonferroni correction. All differences are significant between the empirical kernel output due to the extremely large sample size. All box-and-whisker plots are constructed via the median value as the central line, the interquartile range (IQR) as the box, and the whiskers denoting the minimum and maximum value of the distribution. Outliers are defined as points that lie outside of ±$\pm 1.5\times\text{IQR}$ and were excluded from the plot for clarity, though all points were included in statistical analysis.
We investigate this claim empirically by computing the activations of each kernel size group over the entire test dataset, grouping activation values by whether they were generated by artifact-containing PPG signals or clean signals. Essentially, we perform the convolution and weighting operation for each set of kernels grouped by size (long, moderate, or short), along the entire test dataset, concatenating the output values to one array. We further segregate the output values depending on whether they were generated by PPG signals labeled as artifact or clean. In the framework of our model, this is equivalent to computing:
@@
We find the claims generated by our mathematical analysis of kernel importance to be consistent with our empirical findings, with long kernels producing positive values most of the time, moderate kernels producing mostly negative or mixed positive/negative values (as seen in the medium-sized model), and short kernels producing mostly negative values, as shown in Figure 3. We demonstrate one such 30-second example of this in Figure 4. The agreement between our mathematical exploration of kernel importance and our empirical exploration of the values produced by various kernel groups over the test sets demonstrates that we can draw meaningful conclusions of our model without relying on post hoc explanability methods. These results are shown in Figure 3 by comparing the top and bottom rows.
@@
Next, we note that because our smallest model consists of only 12 kernels, we can easily visually inspect each kernel. They are all shown in Appendix Figure 8. This is in stark contrast compared to most deep-learning models, where the parameters are typically so numerous that it is intractable to manually inspect each parameter; here, all of the kernels can be shown on one page. Furthermore, parameters in deep neural networks are typically uninterpretable, whereas our model represents a template-matching paradigm for waveforms.
@@ 5.1.3 Pruning Experiments
We find that applying our pruning procedure on long kernels provides the greatest parameter reduction with the least performance trade-off. Pruning moderate and/or short kernels impacts performance more heavily and ultimately does not prune as many parameters as pruning long kernels, since each long kernel contains more parameters by definition. Ultimately, the choice of prune-to-performance ratio is up to the user and specific use case, though we provide example pruning schemes in Appendix Table 6 which attempt to maintain performance around 96% of the mean original performance. Notably, as the starting model size increases, we are able to remove progressively more parameters with less of a decrease in performance. We hypothesize this may be due to increased capacity of larger models with greater learned redundancy.
@@ 5.2 Atrial Fibrillation Detection
5.2.1 Classification Performance
We tested our 12-kernel, 72-kernel, and 384-kernel models against a 16-layer ResNet and a previously state-of-the-art convolutional neural network. We train and validate on the computing in cardiology atrial fibrillation detection dataset [49] and test on a holdout set collected by Zheng et al. [50, 51] (training/testing methodology and dataset descriptions are discussed in the Methods section). On the 10-fold cross-validation set, the medium and large-sized SMoLK models outperform or match all other baselines in terms of AUC-ROC, a critical clinical metric (Table 2), despite having dramatically fewer parameters (less than 1% of ResNet-16’s parameter count). On the hold-out test set, all SMoLK models achieve superior AUC-ROC to the other baselines for normal and atrial fibrillation detection (Table 3), though they underperform in detection of unknown “other” rhythms compared to the two highest parameter count models we tested: a 16-layer ResNet (referred to as ResNet-16) and a previously state-of-the-art convolutional neural network (referred to as Jia-CNN1).
@@ 5.2.1 Classification Performance
Jia-CNN1 $0.710\pm 0.029$ $0.856\pm 0.013$ $0.900\pm 0.026$ $0.789\pm 0.013$ $93.6$K
Jia-CNN2 $0.671\pm 0.012$ $0.838\pm 0.010$ $0.914\pm 0.012$ $0.749\pm 0.014$ $28.3$K
Jia-CNN3 $0.720\pm 0.011$ $0.867\pm 0.008$ $0.934\pm 0.012$ $0.786\pm 0.011$ $33.4$K
Table 2: Atrial Fibrillation Detection. Metrics on the train set, computed via 10-fold cross-validation.
@@
We also find that our models train and generalize significantly better in the low-data regime. Training on only 1% of the dataset, our method achieves higher F1 and AUC-ROC scores than all baselines on both the train and holdout set, with ResNet-16 yielding performance consistent with random guessing. We hypothesized this could be due to overparameterization of our ResNet-16 model, thus as a control, we train a simple four-layer convolutional neural network with nearly the same parameter count as our large learned kernel model. We find that even when controlling for parameter count, deep neural networks struggle to learn in the low data regime, achieving near-random performance, while our method maintains a stable performance across parameter counts (Table 4). This is also evident at the varying parameter scales of the Jia-CNN models, though notably, Jia-CNN3 outperforms ResNet-16 in the low-data regime.
@@ 5.2.2 Interpretability
To demonstrate the built-in interpretability of our method for classification tasks, we train a large learned kernel model to distinguish between $1^{\circ}$ atrioventricular block and sinus rhythm on the Zheng et al. dataset. We then apply Algorithm 1 to create a contribution map for each class. We find that our importance maps reliably identify the PR-interval (the defining factor) as the most contributory factor to the classification of $1^{\circ}$ atrioventricular block, despite never explicitly being given this signal (Figure 5). Importantly, this is not a post-hoc methodology (such as SHAP or GradCAM) that seeks to explain a model’s behavior but is not guaranteed to be faithful. Instead, this method numerically computes the precise contribution of each part of the signal to the classification.
Figure 5: Selected ECG Interpretations. Various examples of $1^{\circ}$ atrioventricular block from the Zheng et al. dataset, along with the corresponding contribution maps as computed by Algorithm 1. Redder colors indicate greater contribution to the $1^{\circ}$ atrioventricular block, while bluer colors indicate lesser contribution.
@@ 6 Discussion
We proposed a lightweight architecture for medical signal processing, which maintains performance on par or exceeding deep learning methods, all while remaining interpretable and efficient. We demonstrated the efficacy of our approach using PPG artifact detection and single-lead ECG atrial fibrillation detection as case studies. For PPG artifact detection, our approach achieved near state-of-the-art performance on three benchmark datasets with several orders of magnitude fewer parameters than deep learning methods and showed better generalization to out-of-distribution data, as evident by the markedly improved performance in the most out-of-distribution task, the TROIKA dataset. SMoLK’s simplicity and interpretability allow for the inspection of the model’s inner workings, providing insights into the contribution of each kernel to the overall prediction. Our method demonstrates extremely low compute and memory demands, setting it apart from deep-learning-based approaches. This enables high-quality signal assessment on low-power wearable devices.
@@
Our results demonstrated that SMoLK effectively captures relevant features for identifying good and poor-quality PPG signals, and provides a directly interpretable, lightweight, and performant classifier. Our analysis of kernel importance revealed that larger kernels primarily contribute to the identification of poor-quality signal features, while smaller and medium-sized kernels capture normal PPG features. We demonstrate that our single-lead ECG classifier is directly interpretable, correctly identifying defining features of arrhythmias. Importantly, our work demonstrates that for some complex tasks, simple and interpretable models can perform just as well as deep neural networks and are even preferable, in agreement with studies in other domains [52].
@@
We developed a parameter reduction techniques to improve the compute-to-performance ratio, keeping in mind the low-power limitations of wearable devices. We demonstrated that we can remove over 22% of the parameters in large models and maintain above 98% of the original performance on a PPG segmentation task (Appendix Table 6). Owing to the simplicity and near linearity of our model, we can naïvely quantize the model weights with no impact on performance. This is in contrast to deep neural networks, where the extreme sensitivity of the system can amplify and propagate error due to quantization, leading to performance degradation. Quantization is also often a non-trivial process in deep neural networks as a result of dynamic range issues, again owing to the sensitivity of neural networks to small changes in parameter values. SMoLK suffers from none of the aforementioned complexities.
@@
Our largest model can fit in under 100 KB of memory when cast to float16 and can run in under 75 KB after weight pruning. Our smallest model runs using under 3 KB of memory cast to float16. Computing a raw (i.e., without post-process smoothing) signal score is extremely computationally efficient, with our largest model requiring on the order of 10 MFlops/second of signal processing. This memory-efficient and compute-efficient approach means that designers of low-power, wearable devices can now achieve state-of-the-art PPG signal quality assessment on extremely stringent compute budgets. SMoLK is so compute-efficient that it could feasibly be implemented on microcontrollers, such as consumer pulse oximeters.Additionally, implementation of our architecture is simple and straightforward to adapt to lower-level programming languages, whereas deep learning-based methods require the time-intensive re-implementation of complex operations into lower-level devices. Furthermore, for segmentation tasks, SMoLK’s output only has local dependencies on the input signal; one does not need to wait for chunks of signal to be acquired before running the segmentation algorithm, nor does one need to stitch together multiple prediction windows. Instead, the output of the segmentation algorithm can be computed in real-time as the signal is acquired.
We also highlighted potential limitations in the datasets used for evaluation, as some labeling inaccuracies and discrepancies were observed in the test set samples with the poorest performance. This suggests that the practical maximum achievable performance for PPG artifact detection on these test sets may be limited by the quality of the available ground truth labels. Additionally, while we employed several hold-out sets for the PPG task to demonstrate the generalizability of our method, the datasets suffer from a small number of test subjects, limiting the statistical power of our result; further research is imperative to verify that the success of this methodology holds with a more diverse population. We evaluated our model on real-world noise, employing datasets whose distributions likely model the most common sources of noise (e.g., motion artifacts in the TROIKA dataset), though admittedly it was not possible to assess the robustness of our methodology to all possible noise sources. With all of this in mind, our methodology, and the field as a whole, would benefit from validation on large datasets composed of a large number of subjects under conditions that would generate a diversity of noise distributions. We have previously advocated for a broader effort [53] to tackle this daunting task in future work.
@@
In conclusion, our method provides a robust, interpretable, and efficient approach for medical signal processing, making it competitive with deep learning approaches in performance and yet suitable for deployment on low-power devices.
@@
While our findings are promising, further research is needed to validate our approach in other medical signal-processing applications. The application of our method to other physiological signals and tasks could be explored, as well as the integration of our approach with other wearable devices like consumer pulse-oximeters, smart-rings, smart watches, and other low-profile devices for real-time signal quality assessment. This could significantly advance the field of wearable health technology, enabling continuous, real-time monitoring of biomarkers with high accuracy and minimal computational resources.
7 Data Availability
The data has been made publicly available at https://doi.org/10.5281/zenodo.13117608 [64]. We refer to the original works [46, 47, 48, 49, 51, 50] for the unprocessed data, as well as maintain a copy of the pre-processed data in our GitHub repository [64] (https://github.com/SullyChen/SMoLK. The Computing in Cardiology dataset and Zheng et al. dataset are available on PhysioNet.org (https://physionet.org/content/challenge-2017/1.0.0/ and https://doi.org/10.13026/wgex-er52, respsectively).